40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 12 19:26:08 2012, Rijuparna Chakraborty, Update, , cavitymode scan 7x
    Reply  Thu Sep 13 07:36:41 2012, Steve, Update, General, restore conditions 
Message ID: 7376     Entry time: Wed Sep 12 19:26:08 2012     Reply to this: 7378
Author: Rijuparna Chakraborty 
Type: Update 
Subject: cavitymode scan 

 Summary: Recorded the presence of higher order modes in IMC

What I did: Misaligned the flat mirror MC1 by small amount in both pitch and yaw (it was needed to be done cause at the beginning of the experiment no higher order modes were present)  and scanned the cavity for frequency-range 32MHz to 45MHz.

I found the presence of higher order modes around 36.7MHz (1st order)  and 40.6MHz (2nd order) along with two other strong modes near 35MHz and 42.5MHz.


Attachment 1: P120912_11.32.jpg  440 kB  | Hide | Hide all
Attachment 2: P120912_14.13.jpg  425 kB  | Hide | Hide all
Attachment 3: P120912_14.17.jpg  458 kB  | Hide | Hide all
Attachment 4: P120912_11.25.jpg  431 kB  | Hide | Hide all
Attachment 5: P120912_14.09.jpg  424 kB  | Hide | Hide all
Attachment 6: P120912_14.30.jpg  400 kB  | Hide | Hide all
Attachment 7: P120912_14.34.jpg  424 kB  | Hide | Hide all
ELOG V3.1.3-